An annuity providing a rate of return of 4.8% compounded monthly was purchased for $43,000. The annuity pays $500 at the end of each month. What will be the principal portion of Payment 92? (Round your answer to 2 decimal places.)

Principal   $
Answer #1

$ 470.94

Note:

In Calculation, number of period has been rounded to next full period.

Step-1:Calculation of number of payment of annuity
Number of payment =nper(rate,pmt,pv,fv)
=106
Where,
rate = 4.8%/12 = 0.004
pmt = 500
pv = -43,000
fv = 0
Step-2:Calculation of principal balance at the end of payment 91
Principal balance is the present value of monthly payment.
Principal balance =-pv(rate,nper,pmt,fv)
7,265.34
Where,
rate = 4.8%/12 = 0.004
nper = 106-91 = 15
pmt = 500
fv = 0
Step-3:Calculation of principal balance at the end of payment 92
Principal balance =-pv(rate,nper,pmt,fv)
6,794.40
Where,
rate = 4.8%/12 = 0.004
nper = 106-92 = 14
pmt = 500
fv = 0
Step-4:Calculation of principal portion of payment 92
Principal portion of payment 92 = Principal balance at the end of payment 91 - Principal balance at the end of payment 92
= 7,265.34 - 6,794.40
= 470.94